OpenStudy (anonymous):

I have a few problems where you are given a point and an equation and you have to find an equation that is either parallel or perpendicular. I just don't get it at all...

OpenStudy (anonymous):

parallel : the same slope perpendicular : slope flip and opposite site

OpenStudy (anonymous):

I know, but what am I to do with the points I'm given? Do I plug them in the equation or what? I'm confused.

OpenStudy (anonymous):

give me a example

OpenStudy (anonymous):

Okay. One of them is Find an equation of the line meeting the specified conditions. Write the final answer in slope-intercept form. Containing the point (2,6) and parallel to y=-4x-1

OpenStudy (anonymous):

y=-4x-1 if parallel use the same slope : -4 and creative another line y -6 = -4(x - 2) y -6 = -4x +8 y = -4x +8 +6 y = -4x + 14 now line : y=-4x-1 parallel to line y = -4x + 14

OpenStudy (anonymous):

-4 is slope 14 is y intercept

OpenStudy (anonymous):

how do you get the slope when there's only one given pair of points?

OpenStudy (anonymous):

use the frist equation slope y=-4x-1 slope is -4

OpenStudy (anonymous):

Okay. That makes so much more sense now. Thanks.

OpenStudy (anonymous):

The slope between two points is \(\frac{y_2-y_1}{x_2-x_1}\)

OpenStudy (anonymous):

one given pair use slpoe = 6/2 or 3

OpenStudy (anonymous):

If you have the equation for a line y = mx + b, m is the slope.

OpenStudy (anonymous):

So, let's say I have one that has the points (0,-5) and perpendicular to y=6x + 8. How would I go about solving that?

OpenStudy (anonymous):

What is the slope for that line y=6x+8?

OpenStudy (anonymous):

given what I said before about m?

OpenStudy (anonymous):

y=6x + 8 perpendicular use 6 flip = 1/6 and negative side is : - 1/6 y- (-5) =-1/6( x- 0) y +5 = -1/6x y = -1/6x -5 line: y=6x + 8 is: perpendicular to y = -1/6x -5

OpenStudy (anonymous):

Thanks so much. That has helped a lot.

OpenStudy (anonymous):

you are welcome

OpenStudy (anonymous):

Now, what if I have a problem where it only has y? For example, points (0,-9) and parallel to 4y=8

OpenStudy (anonymous):

Solve for y.

OpenStudy (anonymous):

If 4y=8 then y=?

OpenStudy (anonymous):

I know y=2 but I'm lost as to what to do next

OpenStudy (anonymous):

Ok, well. Going back to the form y=mx + b, what is your m?

OpenStudy (anonymous):

would it be 0? or undefined?

OpenStudy (anonymous):

It would be 0. y = 0x + 2 = 2

OpenStudy (anonymous):

y=2 is a straight horizontal line that goes through the y axis at y=2 and doesn't increase or decrease with a change in x.

OpenStudy (anonymous):

So the slope is 0. Now to find a new line that is parallel means that the new line has the same slope.

OpenStudy (anonymous):

But instead of going through y=2, it goes through y=-9

OpenStudy (anonymous):

I ended up with y=0x -9 but I'm not sure if that's right?

OpenStudy (anonymous):

Yes, now simplify.

OpenStudy (anonymous):

4y=8 y = 2 slope =0 y-9= 0(x-0) y = 9 line 4y=8 parallel to line y = 9

OpenStudy (anonymous):

Should be y=-9

OpenStudy (anonymous):

y=-9 if you want perpendicular to line 4y=8

OpenStudy (anonymous):

No. y=-9 is parallel to the line y=2. They both have a 0 slope. One intersects (0,-9) the other intersects (0,-9)

OpenStudy (anonymous):

One intersects(0,-9) and the other intersects(0,2) rather

OpenStudy (anonymous):

I also got y=-9

OpenStudy (anonymous):

A perpendicular line would have an undefined slope -1/0

OpenStudy (anonymous):

I forgot negative 9 y=-9 is right amswer